OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. S(, 1), T(, 4)

OpenStudy (anonymous):

S(1/2,1) T(1/2,4)

OpenStudy (anonymous):

Its asking for just X, not x+y and therefore I'm lost/frustrated.

OpenStudy (anonymous):

direction vector is vertical, for examle (1,0) so: x-1/2 =0 x=1/2

OpenStudy (anonymous):

Okay could you explain that a wee bit more? I'm still lost ..

OpenStudy (anonymous):

in a standart form ax+by=c (a,b) is a vector which is perpendicular to the line. Since the line you whant is vertical, perpendicular vector could be (1,0). Rest just plug some point in to the equation...

OpenStudy (anonymous):



OpenStudy (anonymous):

but what about y? its ax+by=c Ugh...

OpenStudy (anonymous):

y coordinate of the vector is 0. Remmeber our perpendicular vector is (1,0), so 1(x-p)+0*(y-q)=0 x-p=0

OpenStudy (anonymous):

So for my equation what would it be?? I hate this I dont get what numbers go where or what the entire thing even means.

OpenStudy (anonymous):

equation is x=1/2

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

How'd you get that though.....

OpenStudy (anonymous):

I'm sorry I'm not trying to be thick...

OpenStudy (anonymous):

like i said befor: in a standart form ax+by=c (a,b) is a vector which is perpendicular to the line. Since the line you whant is vertical, perpendicular vector could be (1,0).So when i substitute this in the equation ax+by=c I get: 1x+0*y=c, or x=c

OpenStudy (anonymous):

@natnatwebb

OpenStudy (anonymous):

thank you for your help.